# Latent Causal Probing: A Formal Perspective on Probing with Causal Models of Data

### v1

@@ Abstract
As language models (LMs) deliver increasing performance on a range of NLP tasks, probing classifiers have become an indispensable technique in the effort to better understand their inner workings. A typical setup involves (1) defining an auxiliary task consisting of a dataset of text annotated with labels, then (2) supervising small classifiers to predict the labels from the representations of a pretrained LM as it processed the dataset. A high probing accuracy is interpreted as evidence that the LM has learned to perform the auxiliary task as an unsupervised byproduct of its original pretraining objective. Despite the widespread usage of probes, however, the robust design and analysis of probing experiments remains a challenge. We develop a formal perspective on probing using structural causal models (SCM). Specifically, given an SCM which explains the distribution of tokens observed during training, we frame the central hypothesis as whether the LM has learned to represent the latent variables of the SCM. Empirically, we extend a recent study of LMs in the context of a synthetic grid-world navigation task, where having an exact model of the underlying causal structure allows us to draw strong inferences from the result of probing experiments. Our techniques provide robust empirical evidence for the ability of LMs to learn the latent causal concepts underlying text.
@@ 1 Introduction
The theoretical section of this paper develops a formal perspective on probing using the language of structural causal models (SCM). Specifically, given a causal model which explains the distribution of tokens observed during training, we pose the central hypothesis as determining whether the LM has learned to represent the latent causal variables of the SCM: concepts that explain how the text was generated, but are never directly observed during training. We then introduce probes as a means of empirically testing such hypotheses, by extracting the value of the latent concepts given only the LM representations as input. Our setting naturally captures broader questions about the inductive bias of LMs trained solely on text, and the latent concepts they acquire over the course of training (Q3).
@@ 2.1 Background: structural causal models
where $do(\text{Forecast}=\text{sun})$ is a causal intervention. The intervention can be conceptualized as forcing the weather station to forecast a sunny day independent of the weather, thereby severing the weather-forecast-umbrella pathway. Figure 1(b) depicts the SCM post-intervention.
@@ 2.2 Case study: causal structure in programming languages
Jin & Rinard (2023) propose an experiment to study whether LMs are able to ground a sequence of actions into a sequence of states, having only seen examples of the initial and final state during training. Specifically, they train a 350M parameter Transformer (Vaswani et al., 2017) on a corpus of specification-program pairs using a standard causal language modeling objective. The programs are strings in a grid-world navigation language with 5 actions (move, turn_right, turn_left, put_marker, pick_marker), sampled uniformly between lengths 6 and 10, inclusive. The specifications consist of the initial and final state, which are 8x8 grids. Executing the program navigates a single robot in the initial state to the final state. We refer to Jin & Rinard (2023) for more details about the language.
@@
More generally, given observations generated according to some unknown causal mechanism, a learner could propose various SCMs of the underlying causal structure consistent with the observations, then use these SCMs to inform future predictions. A fundamental challenge in learning causal structure is the problem of latent variable induction, or inferring what latent variables to define candidate SCMs over. In this work, we focus on the causal structure of programming languages, where the underlying causal dynamics are governed by a precise formal semantics, and the latent variables are given by program states. Having formally defined semantics and latent variables enables us interpret the results of our probing experiments in an unambiguous way; we refer the reader to Sloman (2005); Feder et al. (2022) for surveys of causal structure in natural language.
@@ 3 Latent causal probing
Language models are latent causal concept learners.
At a high level, given an SCM that models the training data as the observed variables, we probe the LM for representations of the latent variables of the SCM. Our main insight, as illustrated in Figure 2, is that knowing the latent value of $s_{n-1}$ could help predict the observed value of $p_{n}$; hence, an LM trained to predict $p_{n}$ might eventually induce the existence of the latent variable $s_{n-1}$. More generally, we refer to any latent variable with a causal effect on the distribution of the training data as a latent causal concept.
3.1 Probing for latent variables
@@ 3.2 Causal mediation analysis of probing
To what extent can the auxiliary task performance be attributed to what LM learns from the latent causal concepts in its training data?
@@
Although path-specific effects offer a crisp conceptual framework for isolating the contribution of the LM in probing experiments, actually computing NIE is not straightforward. First, picking a proper baseline $M^{\prime}$ is critical; intuitively, if we pick an inappropriate $M^{\prime}$, then the NIE will measure the difference between $M$ and $M^{\prime}$ in addition to the latent causal concepts hypothetically mediated by the LM representations. Second, measuring the effect requires intervening along the path of interest while holding the other paths constant, i.e., we would need to retrain LM according to the baseline $M^{\prime}$, which would render the technique prohibitively expensive for large pretrained LMs.
Let $acc(M_{0},M_{1})$ denote the (expected) auxiliary task accuracy after the LM is trained using the SCM $M_{0}$ and the probe is calibrated and measured on $M_{1}$. The following result addresses these challenges (proof in Appendix C).
@@
$acc(M^{\prime},M^{\prime})$ $\geq acc(M,M)$ (1)
$acc(M,M^{\prime})$ $\geq acc(M^{\prime},M).$ (2)
@@
Intuitively, $M^{\prime}$ is a valid baseline when measuring $M^{\prime}$ is easier than measuring $M$ under both normal or intervened circumstances. The conclusion then states that, so long as $acc(M,M)-acc(M,M^{\prime})>0$ (which can be evaluated by running probe calibration and measurement twice rather than training the LM twice), there is no bias in which SCM is used to train the LM and which is the baseline: the LM representations always mediate a positive amount of the measured effect.
A positive NIE now also has a rigorous interpretation as the LM having learned latent causal concepts, as some positive amount of causal effect is transferred through the representations of the LM. For instance, a positive mediated measurement for inductive bias implies that
The presence of latent causal concepts in the pretraining data causes the LM to learn representations that generalize to unknown data.
@@ 3.3 Discussion
A significant measurement $acc(M,M)-acc(M,M^{\prime})>0$ is interpreted evidence that the LM encodes causal concepts in its representations. We conclude with some remarks.
@@
Our mediation technique requires knowing “what would the text have been if the underlying dynamics were different?”, which could be difficult (especially in non-synthetic domains). Similarly, for non-causal latent variables, such as part-of-speech, producing a hypothesis class $\mathcal{M}$ with more than one SCM may not be possible: what would the data look like in a counterfactual world in which “dog” is actually an adverb? Unfortunately, our analysis shows that a baseline $M^{\prime}$ which induces a different distribution of text is a necessary precondition, since otherwise $\text{NIE}_{M,M^{\prime}}(\theta_{\text{LM}})$ and $\text{NIE}_{M^{\prime},M}(\theta^{\prime}_{\text{LM}})$ cannot both be positive (as $M$ and $M^{\prime}$ are indistinguishable when used to train the LM). Intuitively, we interpret this result as saying any measurement is inherently biased when the auxiliary task has only one “right” answer.
@@ 4 Experiments
We conduct empirical study of whether an LM, trained from scratch on a corpus of program data, learns the latent causal concepts in the underlying data generation process.
@@ 4.2 Results
Figure 4 plots the main results. For all four measurements (deductive knowledge, inductive bias, deductive bias, and inductive knowledge) and across all three probes (linear, 1-layer MLP, 2-layer MLP), the mediated measurements are significantly positive (above the dashed line at 0%) by the end of training. We thus conclude that a positive fraction of the observed measurements of causal concepts can be attributed to what is learned by the LM’s representations. We also note that the linear probe exhibits the lowest mediated measurements of the 3 probes across all four tasks. This suggests that more complex auxiliary tasks require more complex probes and, more generally, highlights the importance of probing frameworks that can accommodate deeper probes. Section A.2 contains further results and analyses.
@@ 5 Related work
Several prior lines of work apply causal techniques to the interpretability of LMs. These works typically intervene on either the model’s representations (Elazar et al., 2021; Geiger et al., 2021; Meng et al., 2022; Abraham et al., 2022; Li et al., 2022) or the model’s inputs (Kaushik et al., 2020; Vig et al., 2020; Gangal & Hovy, 2020; Amini et al., 2023), and analyze the causal effect on the LM’s outputs. In contrast, we present a formal framework that, conceptually, intervenes on the model’s training data, and measures the causal effect on the LM’s internal representations as a proxy for knowledge. To the best of our knowledge, Elazar et al. (2022) is the only other work that studies the causal relationship between the training data and the LM, presenting a technique for estimating the causal effect of dataset statistics on the factuality of LM’s outputs.
Causal knowledge and reasoning in LMs
A number of benchmarks test for causal knowledge in pretrained LMs by eliciting outputs on causal reasoning tasks; we refer to Yang et al. (2023) for a survey. Recently, however, researchers have raised concerns that performance on such benchmarks may not correspond to true causal reasoning (Zhang et al., 2023; Zečević et al., 2023). As Yang et al. (2023) state, “The issue here is that the LLM does not need to actually reason at all: it can simply access its training dataset, which contains millions of stories about weather and umbrellas, and approximately retrieve a response.” Similarly, Wu et al. (2023) present empirical evidence that pretrained LMs fail to reason on counterfactual task variants, and attribute their apparent reasoning capabilities to recall from the training data. In contrast, we conduct our experiments in a controlled setting where the LM is trained from scratch. Our focus is also on the ability of the LM to represent and generalize the latent causal concepts, rather than knowledge of the causal relationships between the concepts. Finally, the hypothesis that LMs can learn latent causal concepts is also highly related to the position developed by Andreas (2022), who argues that LMs could model properties of agents that are likely to have uttered the language in their training data.
Frameworks for probing
@@
A number of works have proposed frameworks toward a more rigorous understanding of probing. One line takes an information-theoretic view on the information represented by the LM (Zhu & Rudzicz, 2020; Pimentel et al., 2020; Voita & Titov, 2020; Pimentel & Cotterell, 2021). In contrast, we use probes to identify causal effects and inductive biases. Immer et al. (2022) propose an interpretation of probing as quantifying the inductive bias of pretrained representations for downstream tasks, but their framework differs significantly from ours in that the model is understood as a representation-probe pair. In contrast, our approach unifies the LM training and probe calibration procedures under a single causal framework for analysis, yielding formal guarantees for the control of probes. Our analysis also reveals settings in which prior techniques, such as control tasks (Hewitt & Liang, 2019), can yield biased estimates of the intended auxiliary measurement.
@@ 6 Conclusion
This paper presents latent causal probing, a probing framework that studies whether LMs learn latent causal concepts as a byproduct of the language modeling objective. Our framework offers robust tools for interpreting experiment results through the lens of causal analysis, and in particular, rigorously controls for the probe’s contribution in learning the auxiliary task. Experimentally, we extend a previous study of whether Transformers can infer the intermediate states that underlie a sequence of actions. Our results provide strong empirical evidence that LMs can induce latent causal concepts from textual pretraining.
@@ Appendix C Proofs
$\text{NIE}_{M,M^{\prime}}(\theta_{\text{LM}})$ $:=acc(M,M)-acc(M^{\prime},M)$ (3)
$\text{NIE}_{M^{\prime},M}(\theta^{\prime}_{\text{LM}})$ $:=acc(M^{\prime},M^{\prime})-acc(M,M^{\prime}),$ (4)
@@
$acc(M^{\prime},M^{\prime})$ $\geq acc(M,M)$ (5)
$acc(M,M^{\prime})$ $\geq acc(M^{\prime},M).$ (6)
@@
$\text{NIE}_{M,M^{\prime}}(\theta_{\text{LM}})$ $\leq acc(M^{\prime},M^{\prime})-acc(M^{\prime},M)$ (7)
@@
$\text{NIE}_{M,M^{\prime}}(\theta_{\text{LM}})$ $=acc(M,M)-acc(M^{\prime},M)$ (9)
$\geq acc(M,M)-acc(M,M^{\prime}).$ (10)
@@
$acc(M,M)-acc(M,M^{\prime})\leq\text{NIE}_{M,M^{\prime}}(\theta_{\text{LM}})\leq\text{NIE}_{M^{\prime},M}(\theta^{\prime}_{\text{LM}})$ (11)

### v2

@@ Abstract
As language models (LMs) deliver increasing performance on a range of NLP tasks, probing classifiers have become an indispensable technique in the effort to better understand their inner workings. A typical setup involves (1) defining an auxiliary task consisting of a dataset of text annotated with labels, then (2) supervising small classifiers to predict the labels from the representations of a pretrained LM as it processed the dataset. A high probing accuracy is interpreted as evidence that the LM has learned to perform the auxiliary task as an unsupervised byproduct of its original pretraining objective. Despite the widespread usage of probes, however, the robust design and analysis of probing experiments remains a challenge. We develop a formal perspective on probing using structural causal models (SCM). Specifically, given an SCM which explains the distribution of tokens observed during training, we frame the central hypothesis as whether the LM has learned to represent the latent variables of the SCM. Empirically, we extend a recent study of LMs in the context of a synthetic grid-world navigation task, where having an exact model of the underlying causal structure allows us to draw strong inferences from the result of probing experiments. Our techniques provide robust empirical evidence for the ability of LMs to induce the latent concepts underlying text.
@@ 1 Introduction
The theoretical section of this paper develops a formal perspective on probing using the language of structural causal models (SCM). Specifically, given a causal model which explains the distribution of tokens observed during training, we pose the central hypothesis as determining whether the LM has learned to represent the latent variables of the SCM: concepts that explain how the text was generated, but are never directly observed during training. We then introduce probes as a means of empirically testing such hypotheses, by extracting the value of the latent concepts given only the LM representations as input. Our setting naturally captures broader questions about the inductive bias of LMs trained solely on text, and the latent concepts they acquire over the course of training (Q3).
@@ 2.1 Background: structural causal models
where $do(\text{Forecast}=\text{sun})$ is a causal intervention. The intervention can be conceptualized as forcing the weather station to forecast sun regardless of the weather, thereby severing the weather-forecast-umbrella pathway. Figure 1(b) depicts the SCM post-intervention.
@@ 2.2 Case study: causal structure in programming languages
Jin & Rinard (2023) propose an experiment to study whether LMs are able to ground a sequence of actions into a sequence of states, having only seen instances of the initial and final state during training. Specifically, they train a 350M parameter Transformer (Vaswani et al., 2017) on a corpus of specification-program examples using a standard causal language modeling objective. The programs are strings in a grid-world navigation language with 5 actions (move, turn_right, turn_left, put_marker, pick_marker), sampled uniformly between lengths 6 and 10, inclusive. The specifications consist of the initial and final state, which are 8x8 grids. Executing the program navigates a single robot in the initial state to the final state. We refer to Jin & Rinard (2023) for more details about the language.
@@
More generally, given observations generated according to some unknown causal mechanism, a learner could propose various SCMs of the underlying causal mechanism consistent with the observations, then use these SCMs to inform future predictions, an approach known as causal learning (Schölkopf et al., 2021). A major challenge in the foregoing approach is the problem of latent variable induction, or inferring the latent variables over which candidate SCMs are to be defined. In this work, we focus on the causal structure of programming languages, where the underlying causal dynamics are governed by a precise formal semantics, and the latent variables are given by program states. Having formally defined semantics and latent variables enables us interpret the results of our probing experiments in an unambiguous way; we refer the reader to Sloman (2005); Feder et al. (2022) for surveys of causal structure in natural language.
@@ 3 Latent causal probing
Language models are latent concept learners.
At a high level, given an SCM that models the training data as the observed variables, we probe the LM for representations of the latent variables of the SCM. Our main insight, as illustrated in Figure 2, is that knowing the latent value of $s_{n-1}$ could help predict the observed value of $p_{n}$; hence, an LM trained to predict $p_{n}$ might eventually induce the existence of the latent variable $s_{n-1}$.
3.1 Probing for latent concepts
@@ 3.2 Causal mediation analysis of probing
To what extent can the auxiliary task performance be attributed to what LM learns from the latent variables in its training data?
@@
Although path-specific effects offer a crisp conceptual framework for isolating the contribution of the LM in probing experiments, actually computing NIE is not straightforward. First, picking a proper baseline $M^{\prime}$ is critical; intuitively, if we pick an inappropriate $M^{\prime}$, then the NIE will measure the difference between $M$ and $M^{\prime}$ in addition to the latent variables hypothetically mediated by the LM representations. Second, measuring the effect requires intervening along the path of interest while holding the other paths constant, i.e., we would need to retrain LM according to the baseline $M^{\prime}$, which would render the technique prohibitively expensive for large pretrained LMs.
Let $acc_{aux}(M_{0},M_{1})$ denote the (expected) auxiliary task accuracy after the LM is trained using the SCM $M_{0}$ and the probe is calibrated and measured on $M_{1}$. The following result addresses these challenges (proof in Appendix C).
@@
$acc_{aux}(M^{\prime},M^{\prime})$ $\geq acc_{aux}(M,M)$ (1)
$acc_{aux}(M,M^{\prime})$ $\geq acc_{aux}(M^{\prime},M).$ (2)
@@
Intuitively, $M^{\prime}$ is a valid baseline when measuring $M^{\prime}$ is easier than measuring $M$ under both normal or intervened circumstances. The conclusion then states that, so long as $acc_{aux}(M,M)-acc_{aux}(M,M^{\prime})>0$ (which can be evaluated by running probe calibration and measurement twice rather than training the LM twice), there is no bias in which SCM is used to train the LM and which is the baseline: the LM representations always mediate a positive amount of the measured effect.
A positive NIE now also has a rigorous interpretation as the LM having learned latent concepts, as some positive amount of causal effect is transferred through the representations of the LM. For instance, a positive mediated measurement for inductive bias implies that
The presence of latent causal variables in the pretraining data causes the LM to learn representations that generalize to unknown data.
@@ 3.3 Discussion
A significant measurement $acc_{aux}(M,M)-acc_{aux}(M,M^{\prime})>0$ is interpreted evidence that the LM encodes the latent concepts in its representations. We conclude with some remarks.
@@
Our mediation technique requires knowing “what would the text have been if the underlying dynamics were different?”, which could be difficult (especially in non-synthetic domains). Similarly, for non-causal latent variables, such as part-of-speech, producing a hypothesis class $\mathcal{M}$ with more than one SCM may not be possible: what would the data look like in a counterfactual world in which “dog” is actually an adverb? Unfortunately, our analysis suggests that a baseline $M^{\prime}$ which induces a different distribution of text is a necessary precondition, since otherwise $\text{NIE}_{M,M^{\prime}}(\theta_{\text{LM}})$ and $\text{NIE}_{M^{\prime},M}(\theta^{\prime}_{\text{LM}})$ cannot both be positive (as $M$ and $M^{\prime}$ are indistinguishable when used to train the LM). Intuitively, we interpret this result as saying any measurement is inherently biased when the auxiliary task has only one “right” answer.
@@ 4 Experiments
We conduct empirical study of whether an LM, trained from scratch on a corpus of program data, induces the latent concepts in the underlying data generation process.
@@ 4.2 Results
Figure 4 plots the main results. For all four measurements (deductive knowledge, inductive bias, deductive bias, and inductive knowledge) and across all three probes (linear, 1-layer MLP, 2-layer MLP), the mediated measurements are significantly positive (above the dashed line at 0%) by the end of training. We thus conclude that a positive fraction of the observed measurements of latent concepts can be attributed to what is learned by the LM’s representations. We make several additional observations:
Deeper probes are more (generally) more accurate.
The linear probe exhibits the lowest raw and mediated measurements of the 3 probes across all four tasks. Furthermore, the 2-layer MLP exhibits the highest raw measurement in all four tasks, and the highest mediated measurement in two of the four tasks; and only in the case of the deductive bias (Figure 4(f)) does the 1-layer MLP achieve a substantially higher measurement. This suggests that more complex auxiliary tasks require more complex probes and, more generally, highlights the importance of probing frameworks that can robustly accommodate deeper probes.
Raw measurements can be uncorrelated with mediated measurements.
Comparing deductive and inductive knowledge, the raw measurement for inductive knowledge is the substantially higher (inductive knowledge: $\sim$52and $\sim$58vs. deductive knowledge: $\sim$48and $\sim$51%, using a 1-layer and 2-layer MLP, respectively, at the end of training) but the mediated measurements are reversed (inductive knowledge: $\sim$4and $\sim$7vs. deductive knowledge: $\sim$12and $\sim$15%, using a 1-layer and 2-layer MLP, respectively, at the end of training). These results confirm that the raw measurements of probing experiments may be misleading in that a higher raw measurement does not necessarily mean that the LM performs the auxiliary task better.
Section A.2 contains further results and analyses.
@@ 5 Related work
Several prior lines of work apply causal techniques to the interpretability of LMs. These works typically intervene on either the model’s representations (Elazar et al., 2021; Geiger et al., 2021; Meng et al., 2022; Abraham et al., 2022; Li et al., 2022) or the model’s inputs (Kaushik et al., 2020; Vig et al., 2020; Gangal & Hovy, 2020; Amini et al., 2023), and analyze the causal effect on the LM’s outputs. In contrast, we present a formal framework that, conceptually, intervenes on the model’s training data, and measures the causal effect on the LM’s internal representations. To the best of our knowledge, Elazar et al. (2022) is the only other work that studies the causal relationship between the training data and the LM, presenting a technique for estimating the causal effect of dataset statistics on the factuality of LM’s outputs.
World models in LMs.
The hypothesis that LMs can induce latent concepts is related to evidence of world models in LMs, in which the agenda is to understand the extent to which LMs are capable of grounding their inputs to (some representation of) reality. For instance, Abdou et al. (2021) show that LMs’ representations of color terms have geometric structure that reflects human perceptual distance. Li et al. (2021) find that LMs perform entity state tracking over the course of simple stories. Li et al. (2022) present evidence than an LM trained on Othello transcripts develops a representation of the underlying board state. Our work is based on Jin & Rinard (2023), who show that an LM trained only on sequences of instructions and instances of initial and final states in a grid-world navigation task can develop representations of the intermediate states resulting from the sequence of instructions is applied. All these works use some form of supervised classifier trained for an auxiliary objective, but unlike the present work, do not provide any rigorous guarantees that control for the classifier’s ability to fit the auxiliary objective independently of whether the LM has truly learned a world model (or simply represents the text as is). The closest is Jin & Rinard (2023), whose semantic probing interventions follow a similar interventional intuition. However, they do not present any formal analysis, and their baselines also fail to satisfy the requirements of a valid baseline; Appendix B discusses some of the differences in experimental design in greater detail. Finally, our formulation of world models according to the underlying causal mechanism that generated the data is also highly related to the position developed by Andreas (2022), who argues that LMs could model properties of agents that are likely to have uttered the language in their training data.
Frameworks for probing.
@@
A number of works have proposed frameworks toward a more rigorous understanding of probing. One line takes an information-theoretic view on the information represented by the LM (Zhu & Rudzicz, 2020; Pimentel et al., 2020; Voita & Titov, 2020; Pimentel & Cotterell, 2021). Our work uses probes to quantify knowledge and biases in the LM’s learned representations. Immer et al. (2022) propose an interpretation of probing as quantifying the inductive bias of pretrained representations for downstream tasks, but their framework differs significantly from ours in that the model is understood as a representation-probe pair. In contrast, our approach unifies the LM training and probe calibration procedures under a single causal framework for analysis, yielding formal guarantees for the control of probes. Our analysis also reveals settings in which prior techniques, such as control tasks (Hewitt & Liang, 2019), can yield biased estimates of the intended auxiliary measurement.
@@ 6 Conclusion
This paper presents latent causal probing, a probing framework that studies whether LMs induce latent variables as a byproduct of the language modeling objective. Our framework offers robust tools for interpreting experiment results through the lens of causal analysis, and in particular, rigorously controls for the probe’s contribution in learning the auxiliary task. Experimentally, we extend a previous study of whether Transformers can infer the intermediate states that underlie a sequence of actions. Our results provide strong empirical evidence that LMs can induce latent concepts from textual pretraining.
@@ Appendix C Proofs
$\text{NIE}_{M,M^{\prime}}(\theta_{\text{LM}})$ $:=acc_{aux}(M,M)-acc_{aux}(M^{\prime},M)$ (3)
$\text{NIE}_{M^{\prime},M}(\theta^{\prime}_{\text{LM}})$ $:=acc_{aux}(M^{\prime},M^{\prime})-acc_{aux}(M,M^{\prime}),$ (4)
@@
$acc_{aux}(M^{\prime},M^{\prime})$ $\geq acc_{aux}(M,M)$ (5)
$acc_{aux}(M,M^{\prime})$ $\geq acc_{aux}(M^{\prime},M).$ (6)
@@
$\text{NIE}_{M,M^{\prime}}(\theta_{\text{LM}})$ $\leq acc_{aux}(M^{\prime},M^{\prime})-acc_{aux}(M^{\prime},M)$ (7)
@@
$\text{NIE}_{M,M^{\prime}}(\theta_{\text{LM}})$ $=acc_{aux}(M,M)-acc_{aux}(M^{\prime},M)$ (9)
$\geq acc_{aux}(M,M)-acc_{aux}(M,M^{\prime}).$ (10)
@@
$acc_{aux}(M,M)-acc_{aux}(M,M^{\prime})\leq\text{NIE}_{M,M^{\prime}}(\theta_{\text{LM}})\leq\text{NIE}_{M^{\prime},M}(\theta^{\prime}_{\text{LM}})$ (11)